A sample survey on the average total yearly expenditure included 150 students of a certain university. The mean total expenditure per student per year for the sample was 3,000 with a standard deviation of 500. How likely is it that the students spend an average of 3500 per year as claimed by a parent at .01 significant level. 


1
Expert's answer
2021-09-28T15:15:56-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=3500"

"H_1:\mu\not=3500"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.01," "df=n-1=150-1=149" degrees of freedom, and the critical value for a two-tailed test is "t_c = 2.609."

The rejection region for this two-tailed test is "R = \{t: |t| > 2.609\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{3000-3500}{500/\sqrt{150}}=-12.247"

Since it is observed that "|t| = 12.247 >2.609= t_c ," it is then concluded that the null hypothesis is rejected.

Using the P-value approach: The p-value for two-tailes, "\alpha=0.01, df=149," "t=-12.247" is "p\approx0," and since "p = 0 < 0.01=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is different than 3500, at the "\alpha = 0.01" significance level.